3 children sold a total of 468 cups of lemonade at a school's fun-fair. Devi sold thrice as many cups a Fiza. Fiza sold twice as many cups as Ellie. How many cups of lemonade did Ellie sell?
Answer by nerdybill(7384)

3 children sold a total of 468 cups of lemonade at a school's fun-fair. Devi sold thrice as many cups a Fiza. Fiza sold twice as many cups as Ellie. How many cups of lemonade did Ellie sell?
.
Let x = cups Ellie sold
then
2x = cups Fiza sold
3(2x) = 6x = cups Devi sold
.
x + 2x + 6x = 468
9x = 468
x = 52 cups (Ellie)
.
Fiza:
2x = 2(52) = 104 cups
.
Devi:
6x = 6(52) = 312 cups